6- Assume only three digits are used to represent positive integers and also assume the following operations are correct. Determine the base of the numbers shown in each question below. Did any of the additions overflow?

(a) 654 + 013 = 000   (i.e. in what base does the addition of 654 + 013 results in 000, assuming three digit numbers?)

(b) 024 + 043 + 013 + 033 = 223

(c) 024 + 043 + 013 + 033 = 201
Answer #1

a)

654 + 013 = 000

base is 7

yes overflow occur because

654 + 013= 1000

b)024 + 043 + 013 + 033 = 223

here base is 5

no overflow occur

c)  024 + 043 + 013 + 033 = 201

here base is 6

no overflow occur